# Majorana phases beyond neutrinoless double beta decay

### v1

@@ I Introduction
There are several ways to constrain $\sum m_{i}$. Laboratory experiments give bounds of the order of 1 eV [3]. Cosmology gives stronger bounds. Assuming minimal $\Lambda$CDM, which is not a very good description of the data, the DESI fit to BAO + CMB measurements suggests $\sum m_{i}<0.072\,$eV (95% CL) [6]. At the same time, assuming a lower bound on $\sum m_{i}$ as a prior in the DESI analysis, the preference for NO is only mild. We view the mass ordering as an open question and consider values of the lightest neutrino mass up to $0.1$ eV.
@@
where $m_{i}$ are the three physical neutrino masses. The situation is more complicated in trying to rule out the $\nu_{M}{\rm SM}$. It can be excluded in the case of IO, by not observing $0\nu\beta\beta$. However, in case of NO, the $0\nu\beta\beta$ rate may vanish even if neutrinos are Majorana. In the $\nu_{M}{\rm SM}$ and for IO, there is a lower bound from our knowledge of mixing angles and mass-squared differences, $m_{ee}>16\,{\rm meV}$ (at 95% CL) [3]. Using a variety of nuclear matrix element calculations, the current experimental upper bound from $0\nu\beta\beta$ searches is $m_{ee}<(36-156)\,{\rm meV}$ [3]. On the other hand, in the case of NO, for certain values of the lightest neutrino mass, the $0\nu\beta\beta$ rate may be arbitrarily small, even if the neutrino mass terms violate lepton number. In this case, a null result at $0\nu\beta\beta$ experiments will not determine the nature of neutrino masses.
@@
The properties of LNV processes and their dependence on the Majorana phases have been discussed in the literature [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. The prospects for constraining one combination of Majorana phases from $0\nu\beta\beta$ and for establishing $CP$ violation have been studied in Refs. [7, 8, 9, 10]. Ref. [11] studied the role of Majorana phases in manifestly $CP$ violating effects in the lepton sector. Refs. [12, 13, 14] considered the full effective Majorana mass matrix, $m_{\alpha\beta}$, and the experimental bounds on all six elements, and Ref. [15] studied $m_{\alpha\beta}$ in the context of models with extra GeV-scale sterile states. Refs. [16, 17] studied collider signatures of $\Delta L=2$ processes.
While there are ongoing searches for $0\nu\beta\beta$, and thus for probing $m_{ee}$, the prospects of observing other LNV processes are less promising. In the near future, the best sensitivity to a lepton number violating process involving not only electrons will probably come from $\mu^{-}$ to $e^{+}$ conversion. The Mu2e [19] and COMET [20] experiments will search for the $pp\mu^{-}\to nne^{+}$ rate. Mu2e expects to be sensitive at the $10^{-16}$ level [21], many orders of magnitudes away from the much lower rates that could arise in the $\nu_{M}{\rm SM}$ [22, 23], which would be proportional to $m_{\mu e}^{2}$.
@@ III General statements
From Eq. (III) it it clear that, for any $\{\alpha,\beta\}$, $m_{\alpha\beta}^{2}$ depends only on two phases, which can be expressed as pure Majorana phases. This is true for any one $m_{\alpha\beta}$ observable alone. As soon as more than one is considered, there is no choice of basis for the Majorana phases that allows both to be expressed without the use of Dirac phases. Thus, while $m_{\alpha\beta}$ can be expressed as a function of the two Majorana phases in Eq. (17), considering a different $m_{\alpha^{\prime}\beta^{\prime}}$, expressed in the same basis, the phases in the cosines would also include Dirac phases,
@@ VI Conclusion
    The sensitivity of $m_{\alpha\beta}^{2}$ to a specific Majorana phase scales with the product of the corresponding neutrino masses, that is, as the lightest mass goes to zero, the sensitivity to the second Majorana phase is reduced.
@@
    In the regime of a very small lightest neutrino mass, $m_{\ell}$, one Majorana phase, $\Phi_{2o}$, is dominant, and its relation to $m_{\alpha\beta}$ is approximately independent of $m_{\ell}$ itself. The dominance of a single Majorana phase is common to all entries of $m_{\alpha\beta}$, implying that at leading order only one of the six elements is independent in this limit.
@@ Acknowledgements.
ZL thanks Hitoshi Murayama and Yury Kolomensky for related conversations many years ago. The work of YG is supported in part by the NSF grant PHY1316222. The work of ZL is supported in part by the Office of High Energy Physics of the U.S. Department of Energy under contract DE-AC02-05CH11231. The research of SG is supported in part by the U.S. Department of Energy grant number DE-SC0010107. AD is an awardee of the Women’s Postdoctoral Career Development Award.

### v2

@@ I Introduction
There are several ways to constrain $\sum m_{i}$. Laboratory experiments give bounds of the order of 1 eV [3, 6]. Cosmology gives stronger bounds. Assuming minimal $\Lambda$CDM, which is not a very good description of the data, the DESI fit to BAO + CMB measurements suggests $\sum m_{i}<0.072\,$eV (95% CL) [7]. At the same time, assuming a lower bound on $\sum m_{i}$ as a prior in the DESI analysis, the preference for NO is only mild. We view the mass ordering as an open question and consider values of the lightest neutrino mass up to $0.1$ eV.
@@
where $m_{i}$ are the three physical neutrino masses. The situation is more complicated in trying to rule out the $\nu_{M}{\rm SM}$. It can be excluded in the case of IO, by not observing $0\nu\beta\beta$. However, in case of NO, the $0\nu\beta\beta$ rate may vanish even if neutrinos are Majorana. In the $\nu_{M}{\rm SM}$ and for IO, there is a lower bound from our knowledge of mixing angles and mass-squared differences, $m_{ee}>16\,{\rm meV}$ (at 95% CL) [3]. Using a variety of nuclear matrix element calculations, the current experimental upper bound from $0\nu\beta\beta$ searches is $m_{ee}<(22-122)\,{\rm meV}$ [8]. On the other hand, in the case of NO, for certain values of the lightest neutrino mass, the $0\nu\beta\beta$ rate may be arbitrarily small, even if the neutrino mass terms violate lepton number. In this case, a null result at $0\nu\beta\beta$ experiments will not determine the nature of neutrino masses.
@@
The properties of LNV processes and their dependence on the Majorana phases have been discussed in the literature [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. The prospects for constraining one combination of Majorana phases from $0\nu\beta\beta$ and for establishing $CP$ violation have been studied in Refs. [10, 11, 12, 13]. Ref. [14] studied the role of Majorana phases in manifestly $CP$ violating effects in the lepton sector. Refs. [15, 16, 17] considered the full effective Majorana mass matrix, $m_{\alpha\beta}$, and the experimental bounds on all six elements, and Ref. [18] studied $m_{\alpha\beta}$ in the context of models with extra GeV-scale sterile states. Refs. [19, 20] studied collider signatures of $\Delta L=2$ processes.
While there are ongoing searches for $0\nu\beta\beta$, and thus for probing $m_{ee}$, the prospects of observing other LNV processes are less promising. In the near future, the best sensitivity to a lepton number violating process involving not only electrons will probably come from $\mu^{-}$ to $e^{+}$ conversion. The Mu2e [24] and COMET [25] experiments will search for the $pp\mu^{-}\to nne^{+}$ rate. Mu2e expects to be sensitive at the $10^{-16}$ level [26], many orders of magnitudes away from the much lower rates that could arise in the $\nu_{M}{\rm SM}$, of order $(10^{-40})$ [27, 28], which would be proportional to $m_{\mu e}^{2}$.
@@ III General statements
From Eq. (III) it is clear that, for any $\{\alpha,\beta\}$, $m_{\alpha\beta}^{2}$ depends only on two phases, which can be expressed as pure Majorana phases. This is true for any one $m_{\alpha\beta}$ observable alone. As soon as more than one is considered, there is no choice of basis for the Majorana phases that allows both to be expressed without the use of Dirac phases. Thus, while $m_{\alpha\beta}$ can be expressed as a function of the two Majorana phases in Eq. (17), considering a different $m_{\alpha^{\prime}\beta^{\prime}}$, expressed in the same basis, the phases in the cosines would also include Dirac phases,
@@ VI Conclusion
    The sensitivity of $m_{\alpha\beta}^{2}$ to a specific Majorana phase scales with the product of the corresponding neutrino masses, that is, for a small but nonzero lightest mass, $m_{\ell}$, the sensitivity to the second Majorana phase is reduced linearly.
@@
    In the regime of a very small $m_{\ell}$, one Majorana phase, $\Phi_{2o}$, is dominant, and its relation to $m_{\alpha\beta}$ is approximately independent of $m_{\ell}$ itself. The dominance of a single Majorana phase is common to all entries of $m_{\alpha\beta}$, implying that at leading order only one of the six elements is independent in this limit.
@@ Acknowledgements.
ZL thanks Hitoshi Murayama and Yury Kolomensky for related conversations many years ago. The work of YG is supported in part by the NSF grant PHY1316222. The work of ZL is supported in part by the Office of High Energy Physics of the U.S. Department of Energy under contract DE-AC02-05CH11231. The research of SG is supported in part by the U.S. Department of Energy grant number DE-SC0010107. AD is supported by the Women’s postdoctoral career development award of the Weizmann Institute of Science.